OpenStudy (anonymous):

Find the area under the curve y = 1/(7x^3) from x = 1 to x = t and evaluate it for t = 10, t = 100. Then find the total area under this curve for x is greater than or equal to 1.

OpenStudy (ash2326):

Let's plot the function first \[y=\frac 1 {7x^3}\] for \(x\to 0, y\to ?\)

OpenStudy (anonymous):

ok

OpenStudy (ash2326):

what's the value of y, if x approaches to 0

OpenStudy (anonymous):

1/0 so undefined?

OpenStudy (anonymous):

i think intergration will Do it..)

OpenStudy (ash2326):

yes \(y\to \infty\) now when \(x\to \infty\) what value will y approach

OpenStudy (anonymous):

I do not know I need help lol

OpenStudy (anonymous):

wait y should be undefined not infinity?

OpenStudy (ash2326):

when \(x\to \infty\) \[y=\frac 1 \infty=0\] so the function for x>1 will look like this

OpenStudy (ash2326):

yeah it's undefined, because it approaches infinity

OpenStudy (anonymous):

I understand so far

OpenStudy (ash2326):

now let' find the graph. Let's mark x=1 and x=t on the function, t>1